OpenStudy (anonymous):

Using the rational theorem- f(x)=4x^4+13x^3+31x^2+91x+21=

OpenStudy (anonymous):

\[4 x^4+13 x^3+31 x^2+91 x+21=(x+3) (4 x+1) \left(x^2+7\right) \]

OpenStudy (anonymous):

so p=21 and q=4 right?